# Dual Consolidation for Pre-Trained Model-Based Domain-Incremental Learning

### v1

@@ 4.3 Summary of Duct
We summarize the training steps in Algorithm 1. Specifically, facing a new training task, we extract the class centers with $\phi_{0}$. Afterward, we optimize the model and separately consolidate the features and classifiers. The consolidated model ($[W_{o}^{m};W_{n}]^{\top}\phi^{m}_{i}$) is then utilized for testing. We utilize Sinkhorn’s algorithm Sinkhorn & Knopp (1967) to solve the OT problem in Eq. 9.
Discussions on memory cost: During the training process, we need to keep two models in the memory, i.e., the historical consolidated backbone $\phi^{m}_{i}$ and the online backbone $\phi_{i}$ for learning the current task. However, after the learning process of each domain, the current running model is merged into $\phi^{m}_{i}$ (Eq. 5) without needing to be kept in memory. During inference, we utilize the consolidated backbone $\phi^{m}_{i}$ and classifiers $[W_{o}^{m};W_{n}]$ for classification, which is the same as a single backbone. Consequently, Duct can be fairly compared with others.
5 Experiment
(a) Office-Home ViT-B/16 IN1K
(b) DomainNet ViT-B/16 IN1K
(c) CORe50 ViT-B/16 IN1K
(d) CDDB ViT-B/16 IN1K
(e) DomainNet ViT-B/16 IN21K
(f) CORe50 ViT-B/16 IN21K
@@ 5 Experiment
Replay Ratcliff (1990) 84.23$\pm$2.31 83.75$\pm$0.68 64.78$\pm$2.98 61.16$\pm$1.19 85.56$\pm$0.38 92.21$\pm$0.63 66.91$\pm$18.0 63.21$\pm$11.6
iCaRL Rebuffi et al. (2017) 81.66$\pm$2.43 81.11$\pm$1.23 59.89$\pm$2.86 57.46$\pm$2.31 74.43$\pm$3.18 79.86$\pm$2.96 68.43$\pm$18.7 70.50$\pm$16.5
MEMO Zhou et al. (2023) 71.18$\pm$2.76 63.09$\pm$1.80 61.92$\pm$5.39 58.41$\pm$3.20 64.80$\pm$3.16 68.24$\pm$2.41 60.87$\pm$13.4 58.09$\pm$11.7
SimpleCIL Zhou et al. (2024a) 75.69$\pm$5.03 75.72$\pm$0.00 42.95$\pm$4.84 44.08$\pm$0.00 70.92$\pm$0.74 74.80$\pm$0.00 60.80$\pm$4.49 63.40$\pm$0.00
L2P Wang et al. (2022d) 79.72$\pm$4.19 80.03$\pm$1.29 50.45$\pm$4.10 48.72$\pm$2.83 83.57$\pm$0.35 87.87$\pm$0.51 67.33$\pm$5.98 64.45$\pm$5.83
DualPrompt Wang et al. (2022c) 80.20$\pm$3.81 80.85$\pm$0.14 52.28$\pm$3.35 50.46$\pm$3.17 84.53$\pm$0.89 87.27$\pm$1.06 68.33$\pm$7.52 71.41$\pm$1.24
CODA-Prompt Smith et al. (2023) 84.70$\pm$2.94 85.07$\pm$0.34 59.85$\pm$4.49 59.99$\pm$0.88 87.92$\pm$0.41 91.57$\pm$0.69 69.19$\pm$6.11 74.18$\pm$1.33
EASE Zhou et al. (2024b) 81.16$\pm$3.52 76.33$\pm$2.16 50.50$\pm$2.27 43.72$\pm$1.70 86.30$\pm$0.04 87.02$\pm$1.21 67.78$\pm$2.44 64.96$\pm$8.36
RanPAC McDonnell et al. (2023) 82.30$\pm$3.34 82.28$\pm$0.00 55.20$\pm$3.93 54.80$\pm$0.36 79.16$\pm$0.55 81.38$\pm$0.12 78.92$\pm$4.85 80.48$\pm$0.68
S-iPrompt Wang et al. (2022b) 81.50$\pm$3.17 80.51$\pm$0.21 61.16$\pm$4.57 60.46$\pm$0.90 81.95$\pm$0.57 83.38$\pm$0.70 68.51$\pm$7.20 72.76$\pm$0.43
5.1 Experimental Setup
Dataset: Following the benchmark settings Wang et al. (2022c; d; b); Smith et al. (2023) in PTM-based domain-incremental learning, we evaluate the performance on Office-Home Venkateswara et al. (2017), DomainNet Peng et al. (2019), CORe50 Lomonaco & Maltoni (2017), and CDDB-Hard Li et al. (2023). Specifically, Office-Home has four domains, DomainNet has six domains, CORe50 has 11 domains, and CDDB-Hard has five domains. We consider five task orders to shuffle these domains in the DIL setting for a holistic evaluation and report the details in Section B.2.
Comparison methods: In the comparison, we consider two types of methods, including exemplar-based methods, i.e., Replay Ratcliff (1990), iCaRL Rebuffi et al. (2017), MEMO Zhou et al. (2023), and SOTA exemplar-free DIL methods, i.e., SimpleCIL Zhou et al. (2024a), L2P Wang et al. (2022d), DualPrompt Wang et al. (2022c), CODA-Prompt Smith et al. (2023), EASE Zhou et al. (2024b), RanPAC McDonnell et al. (2023), and S-iPrompt Wang et al. (2022b). We utilize the same pre-trained backbone for all compared methods.
Implementation details: We deploy the experiments using PyTorch Paszke et al. (2019) and Pilot Sun et al. (2023) on NVIDIA 4090. Following Wang et al. (2022d); Zhou et al. (2024b), we consider two typical pre-trained weights, i.e., ViT-B/16-IN21K and ViT-B/16-IN1K. Both are pre-trained with ImageNet21K Russakovsky et al. (2015), while the latter is further finetuned on ImageNet1K. We optimize Duct using SGD optimizer with a batch size of 128 for 15 epochs. The learning rate is set to $0.001$. We select 10 exemplars per class for exemplar-based methods using herding Welling (2009) algorithm. In Duct, we set the consolidation parameter $\alpha_{\phi}=0.5,\alpha_{w}=0.5$. The code will be made publicly available upon acceptance.
Performance Measure: Following Wang et al. (2022d; b), we denote the accuracy among all seen domains after learning the $b$-th task as $\mathcal{A}_{b}$, we mainly consider $\mathcal{A}_{B}$ (the performance after learning the last stage) and $\bar{\mathcal{A}}=\frac{1}{B}\sum_{b=1}^{B}\mathcal{A}_{b}$ (the average performance among all incremental stages) for comparison. We also consider the forgetting measure Chaudhry et al. (2018) to measure the relative forgetting degree in DIL.
5.2 Benchmark Comparison
[…]
Besides, we also report the incremental performance among five task orders (average and standard deviation) in Table 1 and Figure 3(a). As we can infer from the table, Duct works robustly among different task orders, showing stable improvements against other state-of-the-art methods.
(a) Multiple trails
(b) Forgetting measure $\downarrow$
(c) Parameter robustness
5.3 Further Analysis
Forgetting Measure: Figure 2 and Table 1 mainly focus on the accuracy measure, which utilizes all seen domains to measure the relative performance. Apart from the accuracy measure, we also follow Wang et al. (2022b; d) to utilize the forgetting measure, which captures the stability of existing knowledge among previous domains. As shown in Figure 3(b), we report the forgetting measure among all compared methods on the CDDB dataset. As we can infer from the figure, several methods suffer from severe forgetting, e.g., Finetune and MEMO, indicating that they are unsuitable for the domain-incremental learning scenario. We also observe that prompt-based methods freeze the backbone to resist feature drift, which tackles the forgetting phenomena. However, Duct still shows the least forgetting (i.e., 0.12) among all compared methods, indicating its strong performance.
Parameter Robustness: There are two major consolidation steps in Duct, i.e., the representation consolidation in Eq. 5 and the classifier consolidation in Eq. 10. These consolidation steps include the merging parameters between a set of models/classifiers, i.e., the backbone merge ratio $\alpha_{\phi}$ and the classifier merge ratio $\alpha_{W}$. In this section, we conduct ablations on the choice of these parameters to investigate the robustness of Duct with change of them. Specifically, we choose $\alpha_{\phi}$ and $\alpha_{W}$ among $\{0.1,0.25,0.5,0.75,1.0\}$, resulting in 25 parameter combinations. We conduct experiments with these parameter combinations on the Office-Home dataset and show the average performance in Figure 3(c). As we can infer from the figure, Duct is generally robust to parameter changes. Besides, since the merge parameters are designed to trade-off between old and new knowledge, we find a small value (e.g., 0.1) or large value (e.g., 1.0) works poorly. The poor performance of these marginal parameters also indicates the importance of the merging process since a small value or large value equals to ignoring part of the important information during consolidation. By contrast, both of them prefer a mild value, i.e., choosing them around {0.25, 0.5} shows better performance. Hence, we suggest setting $\alpha_{\phi}=0.5,\alpha_{w}=0.5$ as the default setting.
Variations $\bar{\mathcal{A}}$ ${\mathcal{A}}_{B}$
Baseline 66.56 63.40
Variation 1 80.36 74.17
Variation 2 81.41 76.82
Variation 3 85.42 80.31
Duct 87.74 82.35
Table 2: Ablation study on differnt modules in Duct.
Figure 4: Before Duct.
Figure 5: After Duct.
Figure 6: Visualizations of the embedding space via t-SNE Van der Maaten & Hinton (2008) on DomainNet. We visualize two incremental stages and utilize dots to represent the first domain and triangles for the second domain. The shadow region denotes the decision boundaries.
Compositional Components: In this section, we conduct experiments to investigate the importance of each module in Duct on CDDB dataset. As shown in Table 2, ‘Baseline’ denotes directly freezing the embedding and extracting class centers as the classifier, which shows poor performance due to the domain gap to the downstream domains. We then utilize Eq. 4 to sequentially merge the backbone to consolidate the representations and denote it as Variation 1. It shows that adding the representation consolidation process drastically improves the performance, indicating the superiority of such unified representation in DIL. However, when switching Eq. 4 to Eq. 5, we find Variation 2 further improves the performance, indicating that task similarity is helpful in obtaining a universal embedding space. Afterward, we combine it with the classifier retraining process in Eq. 8, which is denoted as Variation 3. It shows that the mismatch between consolidated features and the classifiers weakens the performance, and aligning the new classifiers to the embedding helps DIL. When comparing Duct to Variation 3, we find the old classifier transport is also vital to recover former knowledge and resist forgetting, which improves the final accuracy by 2%. Ablation studies verify the efficacy of different modules in Duct.
Visualizations: In this section, we visualize the embedding space to show the effectiveness of Duct on the DomainNet dataset. We consider a two-stage domain-incremental learning scenario, each containing five classes, and utilize t-SNE Van der Maaten & Hinton (2008) to visualize the representation space before and after Duct. We use dots to represent the classes in the first domain and triangles to represent classes in the second domain. We utilize the shadow regions to denote the decision boundary obtained by the classifier weights to represent different classes. As shown in Figure 4, although the embedding space before Duct shows competitive performance, there are two major flaws: 1) There exists the confusion region between yellow and red dots, indicating that previous classes are partially forgotten as data evolves. 2) The embedding space of the same class is still located in different regions, e.g., the purple and green classes, which is not ideal for domain-incremental learning. However, when applying Duct to consolidate the features and classifiers, the above problems are addressed in Figure 5. Specifically, since the consolidated embedding suits all tasks, the forgetting of previous domains is alleviated. Besides, the unified embedding space adequately places the same class of different domains together, favoring the final inference.
6 Conclusion
Domain-incremental learning is a desired ability for real-world learning systems to obtain new knowledge. In this paper, we propose dual consolidation (Duct) to achieve this goal. Since the forgetting in DIL occurs in two aspects, i.e., the embedding and the classifier, we separately design the consolidation technique to handle them. Specifically, we consider merging the pre-trained model with domain-specific task vectors to achieve the unified embedding. To compensate for the mismatch between consolidated features and classifiers, we design the classifier consolidation process by introducing semantic-guided transport. Extensive experiments validate the effectiveness of Duct.
Limitations: Possible limitations include the utilization of PTMs since feature consolidation relies on generalizable initialization. Future work includes extending Duct to non-PTM scenarios.
Appendix
In the main paper, we present a method to prevent forgetting in domain-incremental learning through representation and classifier consolidation. The supplementary material provides additional details on the experimental results mentioned in the main paper, along with extra empirical evaluations and discussions. The organization of the supplementary material is as follows:
  • Section A presents additional experimental results, including the running time comparison, detailed performance on different task orders, and results with other pre-trained weights.
  • Section B introduces the details about the datasets adopted in the main paper, including the number of tasks and images and differently-ordered sequences of tasks.
  • Section C introduces the compared methods adopted in the main paper.
Table 3: Average and last performance of different methods with the 1st task order in Section B.2. The best performance is shown in bold. All methods are implemented with ViT-B/16 IN1K. Methods with $\dagger$ indicate implementations with exemplars (10 per class).
Method Office-Home DomainNet Core50 CDDB
Finetune 73.48 76.23 47.99 35.80 74.44 72.18 50.43 50.23
@@ 6 Conclusion
iCaRL Rebuffi et al. (2017) 77.63 82.03 60.25 59.05 76.79 81.60 49.63 49.77
MEMO Zhou et al. (2023) 70.17 59.80 60.65 60.33 64.89 67.55 53.00 53.67
SimpleCIL Zhou et al. (2024a) 71.60 75.72 39.61 44.08 70.81 74.80 54.52 63.40
L2P Wang et al. (2022d) 74.49 78.44 49.00 51.07 83.47 88.33 58.69 67.89
DualPrompt Wang et al. (2022c) 75.88 80.72 51.92 52.73 85.42 88.09 63.10 71.72
CODA-Prompt Smith et al. (2023) 81.20 85.17 58.49 58.21 87.70 90.85 63.33 72.28
EASE Zhou et al. (2024b) 78.56 77.15 53.04 45.63 86.28 88.98 66.47 72.18
RanPAC McDonnell et al. (2023) 78.72 82.28 53.48 54.97 79.44 81.32 72.47 81.04
S-iPrompt Wang et al. (2022b) 77.32 80.42 58.21 58.88 80.91 82.09 61.75 73.23
@@
Table 4: Average and last performance of different methods with the 2rd task order in Section B.2. The best performance is shown in bold. All methods are implemented with ViT-B/16 IN1K. Methods with $\dagger$ indicate implementations with exemplars (10 per class).
@@
iCaRL Rebuffi et al. (2017) 81.48 81.90 53.75 57.91 75.54 81.17 61.08 76.31
MEMO Zhou et al. (2023) 71.44 62.92 53.41 57.04 61.30 67.52 50.54 48.59
SimpleCIL Zhou et al. (2024a) 68.45 75.72 38.18 44.08 71.63 74.80 63.19 63.40
L2P Wang et al. (2022d) 75.30 79.03 45.15 50.90 83.63 87.57 67.01 70.56
DualPrompt Wang et al. (2022c) 75.41 80.70 46.31 52.29 84.09 87.55 59.85 72.07
CODA-Prompt Smith et al. (2023) 81.21 84.53 52.86 57.38 87.91 91.21 63.05 73.53
EASE Zhou et al. (2024b) 75.96 74.76 47.31 44.62 86.32 86.69 64.08 69.19
RanPAC McDonnell et al. (2023) 78.03 82.28 50.11 54.98 79.44 81.32 75.00 81.04
S-iPrompt Wang et al. (2022b) 78.27 80.81 54.51 60.06 82.09 83.72 61.46 72.43
@@
Table 5: Average and last performance of different methods with the 3rd task order in Section B.2. The best performance is shown in bold. All methods are implemented with ViT-B/16 IN1K. Methods with $\dagger$ indicate implementations with exemplars (10 per class).
@@
iCaRL Rebuffi et al. (2017) 84.40 81.64 61.74 54.41 75.54 81.17 88.66 86.05
MEMO Zhou et al. (2023) 75.50 63.43 62.97 56.99 68.42 71.78 53.94 52.55
SimpleCIL Zhou et al. (2024a) 76.32 75.72 45.74 44.08 69.67 74.80 56.64 63.40
L2P Wang et al. (2022d) 81.94 79.57 52.16 45.05 84.21 88.24 67.87 67.87
DualPrompt Wang et al. (2022c) 82.45 80.79 54.01 49.28 84.91 88.39 66.26 71.19
CODA-Prompt Smith et al. (2023) 63.05 73.53 60.91 56.08 88.58 91.20 68.77 76.26
EASE Zhou et al. (2024b) 81.45 74.76 52.19 40.81 86.26 85.26 67.48 72.18
RanPAC McDonnell et al. (2023) 83.84 82.28 57.57 54.08 78.07 81.62 78.46 81.04
S-iPrompt Wang et al. (2022b) 83.40 80.68 64.02 61.22 82.65 84.20 69.30 72.99
@@
Table 6: Average and last performance of different methods with the 4th task order in Section B.2. The best performance is shown in bold. All methods are implemented with ViT-B/16 IN1K. Methods with $\dagger$ indicate implementations with exemplars (10 per class).
@@
iCaRL Rebuffi et al. (2017) 83.38 78.71 58.30 51.55 68.25 72.91 92.67 88.60
MEMO Zhou et al. (2023) 71.84 65.06 62.28 54.18 68.20 69.76 87.53 80.93
SimpleCIL Zhou et al. (2024a) 81.31 75.72 40.06 44.08 71.73 74.80 66.56 63.40
L2P Wang et al. (2022d) 85.49 81.71 48.59 45.47 83.35 87.01 77.38 61.50
DualPrompt Wang et al. (2022c) 84.74 81.02 50.90 46.77 85.23 86.89 81.46 72.89
CODA-Prompt Smith et al. (2023) 71.06 74.84 60.17 55.89 88.10 91.79 79.73 73.98
EASE Zhou et al. (2024b) 84.69 74.76 48.28 42.92 86.30 86.69 70.23 50.48
RanPAC McDonnell et al. (2023) 86.59 82.28 54.97 53.48 79.31 81.32 85.28 79.66
S-iPrompt Wang et al. (2022b) 85.52 80.19 61.32 60.89 82.11 83.50 81.28 73.06
@@
Table 7: Average and last performance of different methods with the 5th task order in Section B.2. The best performance is shown in bold. All methods are implemented with ViT-B/16 IN1K. Methods with $\dagger$ indicate implemented with exemplars (10 per class).
@@
iCaRL Rebuffi et al. (2017) 80.99 81.26 62.11 54.14 74.86 77.55 50.10 49.77
MEMO Zhou et al. (2023) 66.95 64.24 70.29 63.49 61.18 64.61 53.36 50.35
SimpleCIL Zhou et al. (2024a) 80.75 75.72 51.15 44.08 70.78 74.80 63.06 63.40
L2P Wang et al. (2022d) 81.37 81.38 57.33 51.13 83.21 88.22 65.68 54.42
DualPrompt Wang et al. (2022c) 82.55 81.00 57.33 48.12 82.99 83.48 71.00 69.18
CODA-Prompt Smith et al. (2023) 63.33 72.28 66.85 57.41 87.33 92.81 71.06 74.84
EASE Zhou et al. (2024b) 85.13 80.23 51.68 44.62 86.37 87.47 70.65 60.75
RanPAC McDonnell et al. (2023) 84.31 82.28 61.48 54.98 79.52 81.32 83.39 79.64
S-iPrompt Wang et al. (2022b) 82.70 80.47 67.73 61.23 81.98 83.38 68.76 72.09
@@
Appendix A Supplied Experimental Results
A.1 Running time comparison
A.2 Performance of different task orders
In the main paper, we conduct experiments on the benchmark datasets with five task orders and report the average performance. In this section, we report the performance on each order in Table 3, 4, 5, 6, 7. The task sequences are reported in Section B.2.
Table 8: Details on domain size, train/test split, and instance number of the benchmark datasets. The dataset split and selection follows Wang et al. (2022c; d; b); Smith et al. (2023).
@@ A.2 Performance of different task orders
A.3 Different backbones
In the main paper, we mainly report the performance with ViT-B/16-IN1K. Since Duct is robust with the change of backbones, we report the performance with ViT-B/16-IN21K in this section. Apart from the benchmark datasets reported in Figure 2, the performance on the rest of the benchmarks is shown in Figure 8.
(a) Office-Home
(b) CDDB
Figure 8: Incremental performance of different methods with ViT-B/16 IN21K. We report the performance gap after the last incremental stage between Duct and the runner-up method at the end of the line.
@@ A.3 Different backbones
Appendix B Dataset Details

### v2

@@ Appendix I Algorithm Pipeline
Appendix II Supplied Experimental Results
II.1 Performance of different task orders
In the main paper, we conduct experiments on the benchmark datasets with five task orders and report the average performance. In this section, we report the performance on each order in Table 3, 4, 5, 6, 7. The task sequences are reported in Section III.2.
Table 3: Average and last performance of different methods with the 1st task order in Section III.2. The best performance is shown in bold. All methods are implemented with ViT-B/16 IN1K. Methods with $\dagger$ indicate implementations with exemplars (10 per class).
@@ II.1 Performance of different task orders
Finetune 73.48 76.23 47.99 35.80 74.44 72.18 50.23 50.23
@@
iCaRL [52] 77.63 82.03 60.25 59.05 76.79 81.60 49.63 49.77
MEMO [85] 70.17 59.80 60.65 60.33 64.89 67.55 53.00 53.67
SimpleCIL [88] 71.60 75.72 39.61 44.08 70.81 74.80 54.52 63.40
L2P [70] 74.49 78.44 49.00 51.07 83.47 88.33 58.69 67.89
DualPrompt [69] 75.88 80.72 51.92 52.73 85.42 88.09 63.10 71.72
CODA-Prompt [58] 81.20 85.17 58.49 58.21 87.70 90.85 63.33 72.28
EASE [86] 78.56 77.15 53.04 45.63 86.28 88.98 66.47 72.18
RanPAC [40] 78.72 82.28 53.48 54.97 79.44 81.32 72.47 81.04
S-iPrompt [68] 77.32 80.42 58.21 58.88 80.91 82.09 61.75 73.23
@@
Table 4: Average and last performance of different methods with the 2rd task order in Section III.2. The best performance is shown in bold. All methods are implemented with ViT-B/16 IN1K. Methods with $\dagger$ indicate implementations with exemplars (10 per class).
@@
iCaRL [52] 81.48 81.90 53.75 57.91 75.54 81.17 61.08 76.31
MEMO [85] 71.44 62.92 53.41 57.04 61.30 67.52 50.54 48.59
SimpleCIL [88] 68.45 75.72 38.18 44.08 71.63 74.80 63.19 63.40
L2P [70] 75.30 79.03 45.15 50.90 83.63 87.57 67.01 70.56
DualPrompt [69] 75.41 80.70 46.31 52.29 84.09 87.55 59.85 72.07
CODA-Prompt [58] 81.21 84.53 52.86 57.38 87.91 91.21 63.05 73.53
EASE [86] 75.96 74.76 47.31 44.62 86.32 86.69 64.08 69.19
RanPAC [40] 78.03 82.28 50.11 54.98 79.44 81.32 75.00 81.04
S-iPrompt [68] 78.27 80.81 54.51 60.06 82.09 83.72 61.46 72.23
@@
Table 5: Average and last performance of different methods with the 3rd task order in Section III.2. The best performance is shown in bold. All methods are implemented with ViT-B/16 IN1K. Methods with $\dagger$ indicate implementations with exemplars (10 per class).
@@
iCaRL [52] 84.40 81.64 61.74 54.41 75.54 81.17 88.66 86.05
MEMO [85] 75.50 63.23 62.97 56.99 68.42 71.78 53.94 52.55
SimpleCIL [88] 76.32 75.72 45.74 44.08 69.67 74.80 56.64 63.40
L2P [70] 81.94 79.57 52.16 45.05 84.21 88.24 67.87 67.87
DualPrompt [69] 82.45 80.79 54.01 49.28 84.91 88.39 66.26 71.19
CODA-Prompt [58] 63.05 73.53 60.91 56.08 88.58 91.20 68.77 76.26
EASE [86] 81.45 74.76 52.19 40.81 86.26 85.26 67.48 72.18
RanPAC [40] 83.84 82.28 57.57 54.08 78.07 81.62 78.46 81.04
S-iPrompt [68] 83.40 80.68 64.02 61.22 82.65 84.20 69.30 72.99
@@
Table 6: Average and last performance of different methods with the 4th task order in Section III.2. The best performance is shown in bold. All methods are implemented with ViT-B/16 IN1K. Methods with $\dagger$ indicate implementations with exemplars (10 per class).
@@
iCaRL [52] 83.38 78.71 58.30 51.55 68.25 72.91 92.67 88.60
MEMO [85] 71.84 65.06 62.28 54.18 68.20 69.76 87.53 80.93
SimpleCIL [88] 81.31 75.72 40.06 44.08 71.73 74.80 66.56 63.40
L2P [70] 85.49 81.71 48.59 45.47 83.35 87.01 77.38 61.50
DualPrompt [69] 84.74 81.02 50.90 46.77 85.23 86.89 81.46 72.89
CODA-Prompt [58] 71.06 74.84 60.17 55.89 88.10 91.79 79.73 73.98
EASE [86] 84.69 74.76 48.28 42.92 86.30 86.69 70.23 50.48
RanPAC [40] 86.59 82.28 54.97 53.48 79.31 81.32 85.28 79.66
S-iPrompt [68] 85.52 80.19 61.32 60.89 82.11 83.50 81.28 73.06
@@
Table 7: Average and last performance of different methods with the 5th task order in Section III.2. The best performance is shown in bold. All methods are implemented with ViT-B/16 IN1K. Methods with $\dagger$ indicate implemented with exemplars (10 per class).
@@
iCaRL [52] 80.99 81.26 62.11 54.14 74.86 77.55 50.10 49.77
MEMO [85] 66.95 64.24 70.29 63.49 61.18 64.61 53.36 50.35
SimpleCIL [88] 80.75 75.72 51.15 44.08 70.78 74.80 63.06 63.40
L2P [70] 81.37 81.38 57.33 51.13 83.21 88.22 65.68 54.42
DualPrompt [69] 82.55 81.00 57.33 48.12 82.99 83.48 71.00 69.18
CODA-Prompt [58] 63.33 72.28 66.85 57.41 87.33 92.81 71.06 74.84
EASE [86] 85.13 80.23 51.68 44.62 86.37 87.47 70.65 60.75
RanPAC [40] 84.31 82.28 61.48 54.98 79.52 81.32 83.39 79.64
S-iPrompt [68] 82.70 80.47 67.73 61.23 81.98 83.38 68.76 72.09
@@
Table 8: Details on domain size, train/test split, and instance number of the benchmark datasets. The dataset split and selection follows [69, 70, 68, 58].
@@
II.2 Different backbones
In the main paper, we mainly report the performance using ViT-B/16-IN1K. In this section, we present performance on the remaining benchmark in Figure 7 and all the experimental results on ViT-B/16-IN21K in Figure 8, demonstrating the model’s robustness to changes in backbones.
II.3 Running time comparison
Appendix III Dataset Details